Assume that the team plays 7 games.

If the team is equally likely to win as to lose each game, what is the probability that they win a string of at least 5 games in a row?
Answer #1

Game played, n = 7

Probability of winning = 1/2 = 0.5

Total number of ways = 27 = 128

The possibilities that lead to 5 or more wins in a row in 7 games are as follows:

WWWWWWW
WWWWWWL
WWWWWLW
WWWWWLL
WLWWWWW
LWWWWWW
LWWWWWL
LLWWWWW

Probability that they win a string of at least 5 games in a row = 8/128 = 0.0625